# Connecting the Dots: LLMs can Infer and Verbalize Latent Structure from Disparate Training Data

### v1

@@ D.3 Finetuning hyperparameters
D.4 Training task details
@@ D.5.4 Other evaluations
Appendix E Mixture of Functions task details
E.1 Additional results
E.1.1 Inductive OOCR results
@@
Figure 27: Results for Mixture of Functions on GPT-3.5 vs. GPT-4. “Regression” is the training task. Output 1–Output 3 stand for the three assistant messages that the model has to predict, corresponding to predicting function outputs with 0, 1, and 2 in-context examples (see Section E.4.1 for details). “Reflection” and “Other” are inductive OOCR evaluations.
E.1.2 In-context learning results
@@ E.2 Finetuning hyperparameters
E.3 Training task details
@@ E.4 Evaluation details
E.4.1 Training task evaluations
@@ E.4.3 Other
Appendix F Parity Learning task details
@@ F.2 Evaluation Dataset
F.2.1 Training task evaluation
@@ F.2.2 OOCR evaluation
F.3 Experiment Setup
F.3.1 Measuring Performance
@@ F.4 Results on GPT
Figure 29: GPT models (left: GPT-3.5, right: GPT-4) finetuned to compute the parity of unknown binary variables are able to use the variables in other contexts. Each column corresponds to a type of evaluation (descriptions and examples are in Section F.2).
F.5 Results on Llama3
@@
Figure 30: Llama3 models (left: 8B, right: 70B) finetuned to compute the parity of unknown binary variables are able to use the variables in other contexts. Each column corresponds to a type of evaluation (descriptions and examples are in Section F.2).
F.6 In-Context Results

### v2

@@ E.4 Finetuning hyperparameters
E.5 Training task details
@@ E.6.4 Other evaluations
Appendix F Mixture of Functions task details
F.1 Additional results
F.1.1 Inductive OOCR results
@@
Figure 27: Results for Mixture of Functions on GPT-3.5 vs. GPT-4. “Regression” is the training task. Output 1–Output 3 stand for the three assistant messages that the model has to predict, corresponding to predicting function outputs with 0, 1, and 2 in-context examples (see Section F.4.1 for details). “Reflection” and “Other” are inductive OOCR evaluations.
F.1.2 In-context learning results
@@ F.2 Finetuning hyperparameters
F.3 Training task details
@@ F.4 Evaluation details
F.4.1 Training task evaluations
@@ F.4.3 Other
Appendix G Parity Learning task details
@@ G.2 Evaluation Dataset
G.2.1 Training task evaluation
@@ G.2.2 OOCR evaluation
G.3 Experiment Setup
G.3.1 Measuring Performance
@@ G.4 Results on GPT
Figure 29: GPT models (left: GPT-3.5, right: GPT-4) finetuned to compute the parity of unknown binary variables are able to use the variables in other contexts. Each column corresponds to a type of evaluation (descriptions and examples are in Section G.2).
G.5 Results on Llama3
@@
Figure 30: Llama3 models (left: 8B, right: 70B) finetuned to compute the parity of unknown binary variables are able to use the variables in other contexts. Each column corresponds to a type of evaluation (descriptions and examples are in Section G.2).
G.6 In-Context Results